3) Show that the exponential pdf, {x(x) = () ex12, x>0 has the lack of memory property. That is, show that the conditional
Answer #1

-k -x 8. P(X 2k)= [ f(x)dx = î d : = e 2 -(1) %3D P(X2s+i|X>s)= P[(X2s+tnX2s)] P(X 2s) P(X 2s+t) P(X 2s) -(s+t) (by (1)) -s =